OpenStudy (anonymous):

Solve the equation. Write a reason for each step 5×=4×+8

OpenStudy (anonymous):

bring the 4x over with the 5x so you get 5x-4x=8 then you subtract 5-4=1 and 1 is equivalent to just x so it would be x=8

OpenStudy (gibbs):

Subtract 4x from both sides 5x−4x=8 2 Simplify 5x−4x to x x=8

OpenStudy (anonymous):

Thank you